# arXiv:2603.03056v1[cs.LG] 03 Mar 2026

### v1

@@ 7 Conclusion and further work
The main contribution of this work is an incremental $k$-NN algorithm that guarantees connectedness of the neighborhood graph and its analysis. In contrast to related approaches, the proposed algorithm achieves connectedness using only the search for nearest neighbours without requiring additional information to ensure the graph’s connectivity.
@@
As shown in Section 4, the neighborhood graph created with Algorithm 1 will always be connected. We show (in Section 3) that $k$-NN neighborhood graphs inferred on realistic datasets can exhibit disconnected components for commonly used values of $k$ – as well as up to $k=20$ (see Section 6). Furthermore, when compared to standard $k$-NN, clustering performance is significantly improved for low values of $k$ and of comparable performance for higher values of $k$. Although the incremental nature of the Algorithm 1 makes it sensitive to the ordering of the nodes, we show that the variance of the resulting graphs has little impact on the clustering performance. In Section 6.1.1, we show that using bigger embedding models for the initial high-dimensional embedding expectedly yields consistently better results when used with our approach. Unlike approaches described in Section 2 that show improved clustering performance with the inclusion of MST, the presented incremental approach does not require any global features in order to guarantee connectedness and does not benefit from the addition of global features like MST.
@@
Proposed incremental $k$-NN neighborhood graph construction is simple and effective, but it may be suboptimal in its initial steps. Nodes that are processed at the start of the algorithm have a higher probability of connecting to a large number of weakly related nodes and disproportionally influence the information encoded with a graph. Future work can analyze alternative strategies to approach the early steps of the algorithm and reduce the risk of suboptimal performance. The neighborhood graph produced by the proposed algorithm can be efficiently extended and allows both the addition of new and the deletion of the last added nodes. Those properties were not leveraged in the clustering evaluation as we only evaluated the final clustering performance. Future work can exploit those properties by focusing on applications where new data is continuously added (streaming data) or invalidated. We have shown that in order to get clustering performance comparable to spectral clustering on $k$-NN neighborhood graph, the nearest neighbor relation does not need to be exact. Future work can explore using approximate $k$-NN search, e.g., using the algorithm described in [malkov2018efficient].
One of the open problems in graph-based approaches to clustering is the recalculation of all intermediate steps when new data points are introduced [Mondal2024]. With the incremental construction of the neighborhood graph, we make an initial step in addressing this issue, as the existing neighborhood graph can be reused and easily extended with new data. This property can be readily leveraged in combination with backpropagation-based eigen decomposition [10.5555/3454287.3454571], or a variant of efficient update of eigenvectors [DHANJAL2014440]. Future work should apply those approaches and directly quantify the benefits of an incremental approach to data with a temporal dimension. Such evaluation was done for temporal community detection in [Sattar2023], where authors specifically underline the necessity for new scalable graph algorithms for temporal data – a challenge for which this work provides a readily applicable solution through its approach to incremental neighborhood graph construction.
@@ Acknowledgments
The work was partially supported by the Slovenian Research and Innovation Agency (ARIS) core research programmes P2-0103 and P6-0411, as well as the projects EMMA (L2-50070), and LLM4DH (GC-0002). The work of Boshko Koloski was supported by a young researcher grant PR-12394. The work was also supported by the EU through ERA Chair grant no. 101186647 (AI4DH).

### v2

@@ 7 Conclusion and further work
The main contribution of this work is an incremental $k$-NN algorithm that guarantees connectedness of the neighborhood graph and its analysis. In contrast to related approaches, the proposed algorithm achieves connectedness using only the search for nearest neighbours without requiring additional information to ensure the graph’s connectivity. As shown in Section 4, the neighborhood graph created with Algorithm 1 will always be connected. We show (in Section 3) that $k$-NN neighborhood graphs inferred on realistic datasets can exhibit disconnected components for commonly used values of $k$ – as well as up to $k=20$ (see Section 6). Furthermore, when compared to standard $k$-NN, clustering performance is significantly improved for low values of $k$ and of comparable performance for higher values of $k$. Although the incremental nature of the Algorithm 1 makes it sensitive to the ordering of the nodes, we show that the variance of the resulting graphs has little impact on the clustering performance. In Section 6.2.1, we show that using bigger embedding models for the initial high-dimensional embedding expectedly yields consistently better results when used with our approach. Unlike approaches described in Section 2 that show improved clustering performance with the inclusion of MST, the presented incremental approach does not require any global features in order to guarantee connectedness and does not benefit from the addition of global features like MST.
@@
Proposed incremental $k$-NN neighborhood graph construction is simple and effective, but it may be suboptimal in its initial steps. Nodes that are processed at the start of the algorithm have a higher probability of connecting to a large number of weakly related nodes and disproportionally influence the information encoded with a graph. Future work can analyze alternative strategies to approach the early steps of the algorithm and reduce the risk of suboptimal performance. The neighborhood graph produced by the proposed algorithm can be efficiently extended and allows both the addition of new and the deletion of the last added nodes. Those properties were not leveraged in the clustering evaluation as we only evaluated the final clustering performance. Future work can exploit those properties by focusing on applications where new data is continuously added (streaming data) or invalidated. We have shown that in order to get clustering performance comparable to spectral clustering on $k$-NN neighborhood graph, the nearest neighbor relation does not need to be exact. Future work can explore using approximate $k$-NN search, e.g., using the algorithm described in [malkov2018efficient]. One of the open problems in graph-based approaches to clustering is the recalculation of all intermediate steps when new data points are introduced [Mondal2024]. With the incremental construction of the neighborhood graph, we make an initial step in addressing this issue, as the existing neighborhood graph can be reused and easily extended with new data. This property can be readily leveraged in combination with backpropagation-based eigen decomposition [10.5555/3454287.3454571], or a variant of efficient update of eigenvectors [DHANJAL2014440]. Future work should apply those approaches and directly quantify the benefits of an incremental approach to data with a temporal dimension. Such evaluation was done for temporal community detection in [Sattar2023], where authors specifically underline the necessity for new scalable graph algorithms for temporal data – a challenge for which this work provides a readily applicable solution through its approach to incremental neighborhood graph construction. Availability Code and experimental results available upon acceptnace.
@@ Acknowledgments
The work was partially supported by the Slovenian Research and Innovation Agency (ARIS) core research programmes P2-0103 and P6-0411, as well as the projects EMMA (L2-50070), and LLM4DH (GC-0002). The work of BK was supported by a young researcher grant PR-12394. The work was also supported by the EU through ERA Chair grant no. 101186647 (AI4DH).